# LLAssist: Simple Tools for Automating Literature Review Using Large Language Models

### v1

@@ 1 INTRODUCTION
In response to these mounting challenges, the research community has begun exploring automated solutions to assist in the systematic literature review process (Silva Júnior and Dutra, 2021; Wallace et al., 2010; Tsafnat et al., 2018; Bannach-Brown et al., 2019). Recently, there has been a growing interest in evaluating the potential of Large Language Models (LLMs) for this purpose (Agarwal et al., 2024; Joos et al., 2024; Susnjak, 2023). These advanced AI systems, with their ability to understand and generate human-like text, offer promising avenues for automating various aspects of the literature review process.
A notable contribution to this emerging field comes from Joos et al. (2024), who recently published an extensive evaluation of using LLMs in enhancing the screening process, with results indicating promising potential for reducing human workload. Inspired by these findings, we present LLAssist, a prototype automation tool based on LLM technology. In its current phase, LLAssist is designed as a lightweight and straightforward solution to efficiently process large volumes of articles, extract key information, and assess relevance to specific research questions, thereby streamlining the literature review process. Eventually, we aim LLAssist to be one of the key components for a human-friendly, automated knowledge base and research platform.
By using a predetermined process of key semantics extraction, relevance estimation, and must-read determination, LLAssist provides a simple building block for researchers to expand from the idea. This design choice allows for better adaptability to different research domains and provides researchers with more transparent and interpretable results. It is important to note that while LLAssist shows promise in its current form, it represents an initial step towards more sophisticated literature review automation. The simplicity of its design is intentional, allowing researchers to understand, modify, and build upon its core functionalities as needed for their specific research contexts.
@@ 2.2.3 Automated Analysis
We processed each paper through LLAssist, which performed the following tasks: 1. Extract key semantics (topics, entities, and keywords) from the title and abstract. 2. Evaluate the relevance of each paper to our research questions. 3. Provide relevance and contribution scores (0-1 scale) for each research question. 4. Generate reasoning for relevance and contribution assessments.
@@ 3 TECHNICAL IMPLEMENTATION
The program is designed to work with various LLM providers, including local models (e.g., Ollama Llama 3, Ollama Gemma 2) and cloud-based models (e.g., OpenAI’s GPT-3.5 and GPT-4). This flexibility allows researchers to choose models based on their specific requirements, such as processing speed, accuracy, or data privacy concerns
LLAssist is implemented in C# and utilizes the following key components: 1. Microsoft.SemanticKernel: For interaction with LLM using OpenAI-compatible API, 2. CsvHelper: For reading input and writing output CSV files, and 3. Microsoft.Extensions.Logging: For logging and debugging purposes.
Unlike many existing research tools, LLAssist is developed in C#, a static and strongly typed language, offering benefits such as improved performance, early error detection, and enhanced maintainability (Nanz and Furia, 2015), suitable for further integration into a bigger enterprise system. Additionally, the system will write the CSV output file incrementally to ensure better durability in case of a crash. The sequence diagram can be seen in Figure 1.
@@ 4.1.2 Binary Relevance Decision and Score Distribution
The binary relevance decision and relevance score distribution obtained experiment are shown in Figure 2. Further, the summary of binary relevance and binary contribution decisions are shown in Table 1.
@@ 7 CONCLUSION
In conclusion, LLAssist demonstrated promising capabilities in automating the initial stages of a literature review. Its ability to quickly process and categorize papers offers valuable support to researchers. However, there is room for improvement in utilizing more of the available metadata and fine-tuning the relevance criteria to better differentiate between highly relevant and marginally relevant papers. By leveraging Large Language Models, it offers researchers a valuable tool to efficiently process large volumes of academic literature. While not a replacement for human judgment, LLAssist can significantly reduce the time spent on initial screening and help researchers focus their efforts on the most promising and relevant articles for their research questions, allowing researchers to focus on high-quality work, to achieve higher productivity across expertise and industry.
@@ 8 AVAILABILITY
The source code for LLAssist is freely available at https://github.com/cyharyanto/llassist. We encourage researchers to use, modify, and contribute to this tool to further advance the efficiency of academic literature reviews across various disciplines.
@@ 9 ACKNOWLEDGMENT
We would like to express our deep gratitude to Dr. Arathi Arakala, Dr. Argho Bandyopadhyay, and Dr. Jessica Helmi from RMIT University for their invaluable guidance in systematic literature review methodologies. Their insights into the challenges faced during the screening of papers for surveys and systematic literature reviews shaped the development and focus of LLAssist. We are thankful for their support and shared knowledge.

### v2

@@ 1 INTRODUCTION
In response to these mounting challenges, the research community has begun exploring automated solutions to assist in the systematic literature review process (Silva Júnior and Dutra, 2021; Wallace et al., 2010; Tsafnat et al., 2018; Bannach-Brown et al., 2019). Recently, there has been a growing interest in evaluating the potential of Large Language Models (LLMs) for this purpose (Agarwal et al., 2024; Joos et al., 2024; Susnjak, 2023). These advanced AI systems, with their ability to understand and generate human-like text, offer promising avenues for automating various aspects of the literature review process. Yet, various consumer LLM interfaces such as ChatGPT encountered significant challenges in ensuring reliability and consistency of outputs, adhering to rigorous systematic review methodologies, and addressing ethical concerns related to academic integrity (Cong-Lem et al., 2024).
Without transparent, open-source tools for AI-assisted literature review, researchers face potential limitations in research transparency and replicability that potentially perpetuate biases. These challenges include overlooking critical publications due to time constraints, decreased accuracy from cognitive overload during manual screening, and compromised reproducibility from reliance on closed, proprietary AI systems. Current approaches, while promising, have limitations: Agarwal et al. (2024) focused on generating related work sections but not initial screening; Joos et al. (2024) explored computationally intensive multi-agent approaches; and Susnjak (2023) proposed resource-intensive domain-specific fine-tuned LLMs. LLAssist addresses these issues by providing a simple, transparent alternative that researchers can freely modify and adapt.
A notable contribution to this emerging field comes from Joos et al. (2024), who recently published an extensive evaluation of using LLMs in enhancing the screening process, with results indicating promising potential for reducing human workload. Inspired by these findings, we present LLAssist, a prototype automation tool based on LLM technology. We aim LLAssist to be a building block for an automated knowledge base and research platform, and the open-source nature enables researchers to expand from the idea. This design choice allows adaptation to different research domains with transparent and interpretable results. The simplicity of its design is intentional to allow researchers to understand, modify, and build upon its core functionalities.
Our key contributions include: 1. introducing LLAssist, an open-source tool leveraging LLMs to automate key aspects of the literature review process; 2. demonstrating a novel approach to relevance estimation using LLMs; 3. providing insights into different LLM backends’ performance for literature review tasks; and 4. promoting transparency and reproducibility in AI-assisted literature reviews through open-source development.
@@ 2.2.3 Automated Analysis
We processed each paper through LLAssist, which performed the following tasks: 1. Extract key semantics (topics, entities, and keywords) from the title and abstract. 2. Evaluate the relevance of each paper to our research questions. 3. Provide relevance and contribution scores (0-1 scale) for each research question. 4. Generate reasoning for relevance and contribution assessments. Section 3 details the technical implementation of the automated analysis done by the program.
@@ 3 TECHNICAL IMPLEMENTATION
Chain-of-Thought (CoT) prompting is a technique that enhances the reasoning capabilities of large language models by encouraging step-by-step thinking Wei et al. (2022). In LLAssist, CoT is simulated through two main steps: 1. Extract Key Semantics: The model helps to generate intermediate reasoning steps, expanding on key concepts in the prompt. 2. Filtering: From the expanded reasoning chains, the most consistent and logical path will be selected as part of the self-consistency in the CoT Wang et al. (2022). This approach allows LLAssist to break down complex problems into manageable steps, mimicking human-like reasoning.
LLAssist’s approach to literature review automation shares similarities with recent works but also introduces distinct features. Like the LitLLM system proposed by Agarwal et al. (2024), LLAssist employs a modular pipeline that leverages LLMs for key tasks. However, while LitLLM focuses on generating related work sections, LLAssist prioritizes efficient screening and relevance assessment. Compared to the approach described by Joos et al. (2024), which focuses on employing multiple LLM agents for consensus-based filtering, LLAssist currently utilizes a single LLM for both semantics extraction and relevance assessment. While this may sacrifice some of the robustness gained from multi-agent consensus, it simplifies the implementation and potentially reduces processing time.
The program is designed to work with various LLM providers, including local models (e.g., Ollama Llama 3, Ollama Gemma 2) and cloud-based models (e.g., OpenAI’s GPT-3.5 and GPT-4). This flexibility allows researchers to choose models based on their specific requirements, such as processing speed, accuracy, or data privacy concerns. LLAssist is developed in C#, a static and strongly typed language (Nanz and Furia, 2015), suitable for further integration into a bigger enterprise system. The sequence diagram can be seen in Figure 1.
@@ 4.1.2 Binary Relevance Decision and Score Distribution
The binary relevance decision and relevance score distribution obtained experiment are shown in Figure 2. Further, the summary of binary relevance and binary contribution decisions is shown in Table 1.
@@ 7 CONCLUSION
In conclusion, LLAssist demonstrated promising capabilities in automating the initial stages of a literature review. Its ability to quickly process and categorize papers offers valuable support to researchers. However, there is room for improvement in utilizing more of the available metadata and fine-tuning the relevance criteria to better differentiate between highly relevant and marginally relevant papers. By providing a simple, transparent tool for literature review automation, LLAssist not only enhances research efficiency but also promotes open science principles and methodological freedom in the rapidly evolving landscape of AI-assisted research. While not a replacement for human judgment, LLAssist can significantly reduce the time spent on initial screening and help researchers focus their efforts on the most promising and relevant articles for their research questions, allowing researchers to focus on high-quality work, to achieve higher productivity across expertise and industry.
@@ 8 AVAILABILITY
The source code for LLAssist is freely available at https://github.com/cyharyanto/llassist and is being actively developed to improve usability. We encourage researchers to use, modify, and contribute to this tool to further advance the efficiency of academic literature reviews across various disciplines.
@@ 9 ACKNOWLEDGMENT
We would like to express our deep gratitude to Dr. Arathi Arakala, Dr. Argho Bandyopadhyay, and Dr. Jessica Helmi from RMIT University for their invaluable guidance in systematic literature review methodologies.